OpenStudy (anonymous):

The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

\[d = ut + 0.5at^2 \rightarrow d - ut = 0.5at^2\] \[\frac{d - ut}{0.5t^2} = a \rightarrow a =\frac{d - ut}{0.5t^2}\] \[a =\frac{2(d - ut)}{t^2}\] @HazandBoobear

OpenStudy (anonymous):

@dpasingh thanks